OpenStudy (anonymous):

2x³ - x² + kx + 4 = 0 The multiplication of two of the roots is 1. What's the value of k? a) -4 b) 4 c) 0 d) -8 How can I start solving this?

OpenStudy (ash2326):

We have \[ 2x^3-x^2+kx+4=0\] Let's divide the whole equation by 2 we get \[ x^3-\frac{x^2}{2}+\frac{k}{2} \times x+2=0\] We know that standard cubic equation \[x^3+ax^2+bx+c\] if the roots are x, y , z then \[a=-(x+y+z)\] \[b=(xy+yz+zx)\] \[c=-xyz\] Let the roots of given equation be x, y and z then for the given equation \[-1/2=-(x+y+z)\] \[k/2=(xy+yz+zx)\] \[2=-xyz\] It's given that product of two roots is 1, let xy=1 We have\[2=xyz=>z=2\] We have \[ -1/2= -(x+y+z)=> x+y= -3/2\] We have \[k/2=(xy+yz+zx)\] now we have xy=1, x+y=-3/2 and z=2 \[ k/2= 1+z(x+y)\] so \[k/2= 1+ 2\times \frac{-3}{2}\] \[k/2=1-3\] \[k/2=-2=>k=-4\]

OpenStudy (ash2326):

Melinda did you understand?

OpenStudy (anonymous):

I understood until the standard cubic equation...

OpenStudy (ash2326):

See that's the theory of equations Consider a quadratic equation ax^2+bx+c=0 divide it by a x^2+ b/a x + c/a=0 Let its roots be x and y then x+y=-b/a or -b/a= sum of the roots xy= c/a or c/a= sum of the roots taken 2 at a time Say if we have a, b and c as roots, sum of roots taken 2 at a time will be ab+bc+ca Did you understand till now?

OpenStudy (ash2326):

Melinda???

OpenStudy (anonymous):

Wait a sec pls

OpenStudy (ash2326):

sure!!

OpenStudy (anonymous):

I'm following.

OpenStudy (ash2326):

Now if you have a cubic equation with roots x, y and z say x^3+ax^2+bx+c=0 then a=-(sum of the roots)= -(x+y+x) b= sum of the roots taken 2 at a time= (xy+yz+zx) c= -(sum of the roots taken 3 at a time)= -(xyz)

OpenStudy (anonymous):

That's the first time I'm hearing that. I'm writing it down. Interesting. Thank you so much!

OpenStudy (ash2326):

Melinda this is valid for any type of equation x^4+ax^3+bx^2+cx+d=0 Let its roots be w, x, y and z then a= -(w+x+y+z) sum of the roots b=(wx+wy+wz+xy+xz+yz) sum of the roots taken 2 at a time c=-(wxy+wxz+xyz+ywz) sum of the roots taken 3 at a time d=(wxyz) sum of the roots taken 4 at a time

OpenStudy (ash2326):

Did you understand it?

OpenStudy (anonymous):

Oh I see. Yep, I understand it. I have another exercise which I suppose it's the same process. Check it out: If m, p and mp are the 3 roots from x³ + mx² + mpx + p = 0, the sum of the roots will be... Its the same thing, right?

OpenStudy (ash2326):

yeah m+p+mp=-m

OpenStudy (anonymous):

Thanks! I will try to solve it.

OpenStudy (ash2326):

Welcome :)

OpenStudy (anonymous):

I mean this: 2=−xyz It's given that product of two roots is 1, let xy=1 We have 2=xyz=>z=2

OpenStudy (ash2326):

hey @MelindaR it's because when we have \[x^3+ax^2+bx+c=0\] let the roots be x, y and z then \[ x+y+z=-a\] \[xy+yz+zx=b\] \[xyz=-c\] remember that signs change, beginning with -, then + and so here c=4 Sorry I made a mistake it'd be xyz=-4 so z=-2 sorry @MelindaR

OpenStudy (anonymous):

That's ok. But I've tried with z=-2 and I didn't reach any results from the options...

OpenStudy (ash2326):

Let me check once

OpenStudy (ash2326):

I got k=-8, what did you get?

OpenStudy (ash2326):

We have \[2x^3-x^2+kx+4=0\] Divide the whole equation by 2 \[x^3-\frac{1}{2}x^2+\frac{k}{2} x+2=0\] let the roots be a, b and c. \[ a+b+c=\frac{1}{2}\] \[ab+bc+ca=\frac{k}{2}\] and \[abc=-2\] Let ab =1 we get c=-2 so \[a+b+c=\frac{1}{2}\] c=-2 we get \[a+b=\frac{5}{2}\] now \[ab+bc+ca=\frac{k}{2}\] ab=1 \[1+c(b+a)=\frac{k}{2}\] \[b+a=-\frac{5}{2}\] and c=-2 so \[1-2\times \frac{-5}{2}=\frac{k}{2}\] we get \[1-5=\frac{k}{2}\] we get \[k=-8\]

OpenStudy (anonymous):

Mmm I made a silly mistake in the end. Thx a lot